# Graph neural networks for learning liquid simulations in dynamic scenes containing kinematic objects

### v1

@@ Abstract
Simulating particle dynamics with high fidelity is crucial for solving real-world interaction and control tasks involving liquids in design, graphics, and robotics. Recently, data-driven approaches, particularly those based on graph neural networks (GNNs), have shown progress in tackling such problems. However, these approaches are often limited to learning fluid behavior in static free-fall environments or simple manipulation settings involving primitive objects, often overlooking complex interactions with dynamically moving kinematic rigid bodies. Here, we propose a GNN-based framework designed from the ground up to learn the dynamics of liquids under rigid body interactions and active manipulations, where particles are represented as graph nodes and particle-object collisions are handled using surface representations with the bounding volume hierarchy (BVH) algorithm. This approach enables the network to model complex interactions between liquid particles and intricate surface geometries. Our model accurately captures fluid behavior in dynamic settings and can also function as a simulator in static free-fall environments. Despite being trained on a single-object manipulation task of pouring, our model generalizes effectively to environments with unseen objects and novel manipulation tasks such as stirring and scooping. Finally, we show that the learned dynamics can be leveraged to solve control and manipulation tasks using gradient-based optimization methods.
@@ 1 Introduction
Our physical world is shaped by complex interactions between various forms of matter, from solid objects to flowing liquids. Accurately simulating these phenomena is essential across numerous disciplines, including science and engineering, and becomes increasingly critical as we advance toward a future with autonomous agents capable of reasoning, planning, and interacting with objects in novel environments. Moreover, understanding human intuitive physical reasoning and interactions in everyday tasks in the natural environment also requires the capability of simulating physical scenarios [1, 2]. Traditional physics simulators, relying on hand-crafted features and computationally intensive numerical methods, often struggle to scale to the complexity of real-world scenarios.
@@
Recently, learning-based simulators have offered a viable alternative, demonstrating the ability to learn complex dynamics directly from data [3, 4, 5], improve simulation accuracy, and enable natural integration with gradient-based optimization for solving inverse problems [6, 7]. Graph neural networks (GNNs) [8, 9], in particular, have been successful in recent years in capturing the dynamics of diverse physical systems, including liquids, soft bodies, and rigid bodies [10, 11, 12, 13, 14, 15]. While existing GNN-based simulators have shown impressive results in modeling particle dynamics, significant limitations remain. Many current approaches [15, 10] are restricted to simple collision scenarios with basic geometric shapes, hindering their applicability to real-world rigid body interactions. While recently, Mani et al. [14] improved boundary handling with complex rigid bodies via Signed Distance Fields (SDFs), the model is still constrained to free-fall simulations, where the rigid bodies remain fixed in place during simulation. Accurately modeling dynamic liquid interactions with moving objects(kinematic rigid bodies) is essential for applications in sequential decision-making tasks such as robotic control. A simulator designed for free-fall settings can address only design-oriented tasks (e.g., directing fluid flow [16]), and is insufficient for tasks requiring manipulation and control. Crucially, even with highly accurate collision handling, a static framework does not guarantee correct behavior when objects move, as dynamic interactions introduce additional complexities in fluid motion and response. Therefore, achieving accurate dynamic simulations requires careful design choices, including the appropriate GNN input representation and architectural modifications.
@@
To address these challenges, we propose a unified GNN-based framework that simultaneously models complex fluid dynamics and handles intricate interactions of particles with dynamically moving rigid bodies. Our architecture leverages a multi-graph representation where liquid particles and rigid bodies are represented as distinct node sets. Liquid interactions are captured using a distance-based neighborhood criterion, while collisions with rigid bodies are handled via a robust, surface-based collision detection mechanism. We train our model on ground truth simulations of generic pouring and sliding tasks, collectively covering a broad range of liquid-object interactions under dynamic control. Our results show: 1) improved performance both qualitatively and quantitatively over existing GNN-based particle simulators; 2) strong generalisation to unseen object geometries and scene configurations; and 3) successful zero-shot transfer to novel manipulation tasks such as stirring and scooping. Finally, we demonstrate the utility of our model by integrating it into a gradient-based optimization pipeline to solve a pouring task using a Model Predictive Control (MPC) controller. Taken together, these results demonstrate that our proposed model extends the capabilities of GNN-based simulators to include complex interactions of liquids with kinematic rigid bodies.
@@ 2 Related Work
With the rise of deep learning and the increasing preference for end-to-end models, differentiable simulators [17] have become a compelling alternative to traditional physics-based simulators such as Bullet [18] and Mujoco [19] for simulating physics. Unlike conventional simulators, these analytical differentiable simulators enable automatic differentiation, allowing seamless integration with gradient-based optimization techniques [20, 21, 22, 23]. Several frameworks have been developed for fluid dynamics, such as PhiFlow [24], which is framework-agnostic, JAX-Fluids [25], built on JAX, and FluidLab [26], which leverages Taichi [20]. Hybrid approaches have also emerged, combining analytical equations with neural networks to improve accuracy and generalization [27, 28, 29].
@@
Purely learning-based simulators, in particular, Graph Neural Networks (GNNs) [10, 15, 11, 30, 12, 13, 31, 32, 33, 34, 14] have gained significant traction due to their ability to model complex physical interactions. This success can be attributed to the inherent nature of many physical phenomena – the dynamics of rigid bodies, deformable bodies, fluids, and even their coupled interactions can often be described by local interactions between constituent entities. GNNs, with their strong relational inductive bias [9], provide a natural and effective framework for encoding this locality, allowing the model to learn the underlying physics from data. The seminal work by Sanchez-Gonzalez et al. [10] established several key architectural and representational choices that have influenced a number of subsequent works. These include using relative geometric features to define node and edge attributes, employing one-step prediction training instead of full rollout-based supervision, and utilizing message-passing GNNs [35, 36] as the core model. While this work primarily focused on particle-based fluid simulation, later studies successfully extended similar architectures to soft-body [11] and rigid-body [12, 30, 32, 33, 34] simulations.
@@
In physical simulation, accurately modeling collisions is essential. In GNN-based simulators, this is typically addressed through node-based neighborhood interactions. This approach has proven effective for simulating liquid-liquid [10] and deformable body [11] interactions. However, it struggles when dealing with rigid-body collisions, as mesh vertices are often sparse and fail to capture fine-grained object geometry [12, 13, 14]. To address these limitations, Allen et al. [12] proposed handling collisions in the mesh-face space instead of the vertex space, leading to more robust rigid-body interactions. Another widely adopted approach is the use of SDFs for collision handling. SDFs implicitly represent surface geometry by encoding the distance of any point in space to the closest surface on a mesh. This allows for efficient collision detection by evaluating the SDF at query points. Rubanova et al. [13] used learned SDFs to simulate rigid body dynamics while Mani et al. [14] used SDF formulation to learn fluid dynamics in a static free-fall environment, representing all rigid surfaces in the environment as a single union of SDFs. These methods significantly reduce the computational overhead compared to node-based collision handling, offering a scalable alternative for complex physical interactions.
@@
So far, the aforementioned GNN-based simulators have either been designed for static free-fall environments, restricting their use to design optimization tasks [16, 14], or have supported manipulation only in limited settings, typically involving simple rigid bodies like cuboids [15, 10] or spheres [11]. In contrast, our model extends beyond these constraints by generalizing to arbitrary scenes and complex manipulation tasks, marking a significant step toward a fully learned, general-purpose differentiable simulator capable of handling diverse physical interactions.
@@ 3 Method
Our graph network follows the Encode-Process-Decode architecture introduced in [10]. The state of the world is first encoded into a graph where node features are derived from the current state, and edge connections are determined based on factors such as spatial proximity for potential collisions and the underlying geometry of objects. Subsequently, both node and edge features are embedded into a latent space via MLPs. The final encoded graph is passed through $P$ message-passing layers to generate a sequence of $P$ latent graphs. Finally, a decoder, which is also implemented as an MLP, extracts the updated dynamics from task-relevant nodes of the final latent graph.
@@ 3.1 Graph Network
Graph Connectivity in our model is handled using four sets of edges. Following [12, 13], we connect bi-directional edges between object node $o_{i}$ and its corresponding surface nodes $\{m_{ik}\}$, forming edge sets $\mathcal{E}^{OM}$ (object-to-mesh) and $\mathcal{E}^{MO}$ (mesh-to-object). This facilitates the propagation of object motion to the particles through the surface nodes. Interaction between liquid particles is captured using the edge set $\mathcal{E}^{L}$. We connect edges between liquid particles if the distance between them is less than the liquid-liquid connectivity radius $r_{l}$. This promotes local interactions of particles, and we choose radius $r_{l}$ to ensure each particle has roughly 10-20 neighbors.
Crucially, to model the interactions between fluid particles and rigid bodies, we introduce the $\mathcal{E}^{OL}$ edge set. Similar to liquid-liquid interactions, the dynamics of fluid particles are primarily affected when they are near a rigid body’s surface. Notably, this interaction can occur at any arbitrary point on the object’s surface, not just at mesh vertices. To achieve this, we define a function $\mathbf{C}(p^{\prime}_{j},o_{i})$ based on the Boundary Vector Hierarchy (BVH) algorithm [37]. This function takes the mesh geometry of object $o_{i}$ and a query particle $p^{\prime}_{j}$ as input, and efficiently computes the closest point $c^{\prime}_{ij}$ on the surface of the mesh to the query particle. Since this function operates on positions in the object’s local frame, we first transform the liquid particle position $p_{j}$ into this frame using the inverse transformation $\mathcal{T}_{t}^{-1}(p_{j})$ based on the object’s current transformation $\mathcal{T}_{t}$. The closest point is then transformed back to the world coordinate frame:
@@
If the distance $d$ between the particle $p_{j}$ and the transformed closest point $c_{ij}$ is less than the liquid-object connectivity radius, $r_{ol}$, we connect an edge between the particle node and the object node. This approach is similar to [13], where learned SDFs for collision detection between rigid bodies were used. We use a different connectivity radius $r_{ol}$ for liquid-object interactions since our ground truth simulator maintains some separation distance between the object surface and the particles to enforce its boundary condition.
Node features Following the approach of [10, 11, 12, 13, 14], we use a history of velocity information as the main node feature. Specifically, we use the finite-difference estimates of velocity from the previous five timesteps. We calculate these velocity features differently depending on the node type. For liquid particle nodes $\mathcal{V}^{L}$ and mesh surface nodes $\mathcal{V}^{M}$, we calculate them based on their positions over the past five timesteps. For object nodes $\mathcal{V}^{O}$, which represent the rigid bodies, we derive velocity information from the object’s 6-DOF pose at each timestep. This 6-DOF pose information is crucial for the network to learn the rotational dynamics of the objects, in addition to their translational motion. Additionally, we also add features to distinguish between different types of object nodes $\mathcal{V}^{O}$. Our simulation includes kinematic objects (which are controlled) and stationary objects (which remain fixed). To represent this, we use a one-hot encoding to indicate its type. This one-hot encoding is also extended to the mesh nodes $\mathcal{V}^{M}$ since each mesh node belongs to a specific object.
@@
Edge features For the edge sets $\mathcal{E}^{L}$ (liquid-liquid), $\mathcal{E}^{MO}$ (mesh-to-object), and $\mathcal{E}^{OM}$ (object-to-mesh), we adopt the relative encoding technique [10]. The edge features consist of the relative displacement vector between two connected nodes and the scalar distance between them. The calculation of these relative features differs slightly depending on the edge type. For $\mathcal{E}^{L}$, the relative position and distance are computed between the positions of the two connected liquid particles. For $\mathcal{E}^{MO}$ and $\mathcal{E}^{OM}$ edges, we calculate them based on the mesh node position and the center of mass position of the corresponding object, derived from the object node’s 6-DOF pose. For the liquid-object edge set $\mathcal{E}^{OL}$, which is critical for modeling particle collisions with surfaces, we use a specialized feature set. For an edge connected between particle $p_{j}$ and object $o_{i}$, the feature vector is defined as $e_{ij}^{ol}=[c_{ij}-p_{j},c_{ij}-p_{o_{i}},\lVert c_{ij}-p_{j}\rVert,\lVert c_{ij}-p_{o_{i}}\rVert]$. Here, $c_{ij}$ represents the closest point on the surface of object $o_{i}$ to the particle $p_{j}$, and $p_{o_{i}}$ represents the center of mass of object $o_{i}$. These features together give information about the spatial relationship between the connected particle and object, using the closest point on the object’s surface $c_{ij}$, as a reference. The resulting input graph is encoded into separate MLPs for each node type and edge type before passing to the processor block.

### v2

# Learning Particle Dynamics Subject to Rigid Body Manipulations Using Graph Neural Networks

@@ Abstract
Simulating particle dynamics with high fidelity is crucial for solving real-world interaction and control tasks involving liquids in design, graphics, and robotics. Recently, data-driven approaches, particularly those based on graph neural networks (GNNs), have shown progress in tackling such problems. However, these approaches are often limited to learning fluid behavior in static free-fall environments or simple manipulation settings involving primitive objects, often overlooking complex interactions with dynamically moving kinematic rigid bodies. Here, we propose a GNN-based framework designed from the ground up to learn the dynamics of liquids under rigid body interactions and active manipulations, where particles are represented as graph nodes and particle-object collisions are handled using surface representations with the bounding volume hierarchy (BVH) algorithm. Our approach accurately captures fluid behavior in dynamic settings and can also function as a simulator in static free-fall environments. Despite being trained on single-object manipulation tasks, our model generalizes effectively to environments with novel objects and novel manipulation tasks. Finally, we show that the learned dynamics can be leveraged to solve control and manipulation tasks using gradient-based optimization methods.
@@ 1 Introduction
Our physical world is shaped by complex interactions between various forms of matter, from solid objects to flowing liquids. Accurately simulating these phenomena is essential across numerous disciplines, including science and engineering, and becomes increasingly critical as we advance toward a future with autonomous agents capable of reasoning, planning, and interacting with objects in novel environments. Moreover, understanding human intuitive physical reasoning and interactions in everyday tasks in the natural environment also requires the capability of simulating physical scenarios [intuitive_liq, intuitive_phy]. Traditional physics simulators, relying on hand-crafted features and computationally intensive numerical methods, often struggle to scale to the complexity of real-world scenarios.
@@
Recently, learning-based simulators have offered a viable alternative, demonstrating the ability to learn complex dynamics directly from data [he2019learning, wiewel2019latent, ladicky2015data], improve simulation accuracy, and enable natural integration with gradient-based optimization for solving inverse problems [challapalli2021inverse, lin2022diffskill]. Graph neural networks (GNNs) [gnn_theory_2, gnn_theory], in particular, have been successful in recent years in capturing the dynamics of diverse physical systems, including liquids, soft bodies, and rigid bodies [sanchez2020learning, meshgraphnet, fignet, sdf_learning, SURFsUP, dpinet]. While existing GNN-based simulators have shown impressive results in modeling particle dynamics, significant limitations remain. Many current approaches [dpinet, sanchez2020learning] are restricted to simple collision scenarios with basic geometric shapes, hindering their applicability to real-world rigid body interactions. While recently, SURFsUP improved boundary handling with complex rigid bodies via Signed Distance Fields (SDFs), the model is still constrained to free-fall simulations, where the rigid bodies remain fixed in place during simulation. Accurately modeling dynamic liquid interactions with moving objects (kinematic rigid bodies) is essential for applications in sequential decision-making tasks such as robotic control. A simulator designed for free-fall settings can address only design-oriented tasks (e.g., directing fluid flow [inverse_design]), and is insufficient for tasks requiring manipulation and control. Crucially, even with highly accurate collision handling, a static framework does not guarantee correct behavior when objects move, as dynamic interactions introduce additional complexities in fluid motion and response. Therefore, achieving accurate dynamic simulations requires careful design choices, including the appropriate GNN input representation and architectural modifications.
@@
To address these challenges, we propose a unified GNN-based framework that models complex fluid dynamics with dynamically moving rigid bodies. Our architecture leverages a multi-graph representation where liquid particles and rigid bodies are represented as distinct node sets. Liquid–liquid interactions are determined by spatial proximity (particles are connected if they lie within a fixed radius), while liquid–object interactions are determined by particle–surface proximity computed efficiently using the BVH algorithm. We train our model on ground truth simulations of generic pouring and sliding tasks, collectively covering a broad range of liquid-object interactions under dynamic control. Our results show: 1) improved performance both qualitatively and quantitatively over existing GNN-based particle simulators; 2) strong generalisation to unseen object geometries and scene configurations; and 3) successful zero-shot transfer to novel manipulation tasks such as stirring and scooping. Finally, we demonstrate the utility of our model by integrating it into a gradient-based optimization pipeline to solve a pouring task using a Model Predictive Control (MPC) controller. Taken together, these results demonstrate that our proposed model extends the capabilities of GNN-based simulators to include complex interactions of liquids with kinematic rigid bodies.
@@ 2 Related Work
With the rise of deep learning and the increasing preference for end-to-end models, differentiable simulators [diff_review] have become a compelling alternative to traditional physics-based simulators such as Bullet [bullet] and Mujoco [todorov2012mujoco] for simulating physics. Unlike conventional simulators, these analytical differentiable simulators enable automatic differentiation, allowing seamless integration with gradient-based optimization techniques [taichi, brax, JaxMD, genesis]. Several frameworks have been developed for fluid dynamics, such as PhiFlow [PhiFlow], which is framework-agnostic, JAX-Fluids [JaxFluid], built on JAX, and FluidLab [fluidlab], which leverages Taichi [taichi]. Hybrid approaches have also emerged, combining analytical equations with neural networks to improve accuracy and generalization [spnet, cnn_liq, heart_neuralfluid].
@@
Purely learning-based simulators, in particular, Graph Neural Networks (GNNs) [sanchez2020learning, dpinet, meshgraphnet, rigid_discontinuos, fignet, sdf_learning, hopper_control_gnn, gmn, egnn, sgnn, SURFsUP] have gained significant traction due to their ability to model complex physical interactions. This success can be attributed to the inherent nature of many physical phenomena – the dynamics of rigid bodies, deformable bodies, fluids, and even their coupled interactions can often be described by local interactions between constituent entities. GNNs, with their strong relational inductive bias [gnn_theory], provide a natural and effective framework for encoding this locality, allowing the model to learn the underlying physics from data. The seminal work by sanchez2020learning established several key architectural and representational choices that have influenced a number of subsequent works. These include using relative geometric features to define node and edge attributes, employing one-step prediction training instead of full rollout-based supervision, and utilizing message-passing GNNs [messagePassingNN, interaction_network] as the core model. While this work primarily focused on particle-based fluid simulation, later studies successfully extended similar architectures to soft-body [meshgraphnet] and rigid-body [fignet, rigid_discontinuos, gmn, egnn, sgnn] simulations.
@@
In physical simulation, accurately modeling collisions is essential. In GNN-based simulators, this is typically addressed through node-based neighborhood interactions. This approach has proven effective for simulating liquid-liquid [sanchez2020learning] and deformable body [meshgraphnet] interactions. However, it struggles when dealing with rigid-body collisions, as mesh vertices are often sparse and fail to capture fine-grained object geometry [fignet, sdf_learning, SURFsUP]. To address these limitations, fignet proposed handling collisions in the mesh-face space instead of the vertex space, leading to more robust rigid-body interactions. Another widely adopted approach is the use of SDFs for collision handling. SDFs implicitly represent surface geometry by encoding the distance of any point in space to the closest surface on a mesh. This allows for efficient collision detection by evaluating the SDF at query points. sdf_learning used learned SDFs to simulate rigid body dynamics while SURFsUP used SDF formulation to learn fluid dynamics in a static free-fall environment, representing all rigid surfaces in the environment as a single union of SDFs. These methods significantly reduce the computational overhead compared to node-based collision handling, offering a scalable alternative for complex physical interactions.
@@
So far, the aforementioned GNN-based simulators have either been designed for static free-fall environments, restricting their use to design optimization tasks [inverse_design, SURFsUP], or have supported manipulation only in limited settings, typically involving simple rigid bodies like cuboids [dpinet, sanchez2020learning] or spheres [meshgraphnet]. In contrast, our model extends beyond these constraints by generalizing to arbitrary scenes and complex manipulation tasks, marking a significant step toward a fully learned, general-purpose differentiable simulator capable of handling diverse physical interactions.
@@ 3 Method
Our graph network follows the Encode-Process-Decode architecture introduced in [sanchez2020learning]. The state of the world is first encoded into a graph where node features are derived from the current state, and edge connections are determined based on factors such as spatial proximity for potential collisions and the underlying geometry of objects. Subsequently, both node and edge features are embedded into a latent space via MLPs. The final encoded graph is passed through $P$ message-passing layers to generate a sequence of $P$ latent graphs. Finally, a decoder, which is also implemented as an MLP, extracts the updated dynamics from task-relevant nodes of the final latent graph.
@@ 3.1 Graph Network
Graph Connectivity in our model is handled using four sets of edges. Following [fignet, sdf_learning], we connect bi-directional edges between object node $o_{i}$ and its corresponding surface nodes $\{m_{ik}\}$, forming edge sets $\mathcal{E}^{OM}$ (object-to-mesh) and $\mathcal{E}^{MO}$ (mesh-to-object). This facilitates the propagation of object motion to the particles through the surface nodes. Interaction between liquid particles is captured using the edge set $\mathcal{E}^{L}$. We connect edges between liquid particles if the distance between them is less than the liquid-liquid connectivity radius $r_{l}$. This promotes local interactions of particles, and we choose radius $r_{l}$ to ensure each particle has roughly 10-20 neighbors.
Crucially, to model the interactions between fluid particles and rigid bodies, we introduce the $\mathcal{E}^{OL}$ edge set. Similar to liquid-liquid interactions, the dynamics of fluid particles are primarily affected when they are near a rigid body’s surface. Notably, this interaction can occur at any arbitrary point on the object’s surface, not just at mesh vertices. To achieve this, we define a function $\mathbf{C}(p^{\prime}_{j},o_{i})$ based on the Boundary Vector Hierarchy (BVH) algorithm [bvh]. This function takes the mesh geometry of object $o_{i}$ and a query particle $p^{\prime}_{j}$ as input, and efficiently computes the closest point $c^{\prime}_{ij}$ on the surface of the mesh to the query particle. Since this function operates on positions in the object’s local frame, we first transform the liquid particle position $p_{j}$ into this frame using the inverse transformation $\mathcal{T}_{t}^{-1}(p_{j})$ based on the object’s current transformation $\mathcal{T}_{t}$. The closest point is then transformed back to the world coordinate frame:
@@
If the distance $d$ between the particle $p_{j}$ and the transformed closest point $c_{ij}$ is less than the liquid-object connectivity radius, $r_{ol}$, we connect an edge between the particle node and the object node. This approach is similar to [sdf_learning], where learned SDFs for collision detection between rigid bodies were used. We use a different connectivity radius $r_{ol}$ for liquid-object interactions since our ground truth simulator maintains some separation distance between the object surface and the particles to enforce its boundary condition.
Node features Following the approach of [sanchez2020learning, meshgraphnet, fignet, sdf_learning, SURFsUP], we use a history of velocity information as the main node feature. Specifically, we use the finite-difference estimates of velocity from the previous five timesteps. We calculate these velocity features differently depending on the node type. For liquid particle nodes $\mathcal{V}^{L}$ and mesh surface nodes $\mathcal{V}^{M}$, we calculate them based on their positions over the past five timesteps. For object nodes $\mathcal{V}^{O}$, which represent the rigid bodies, we derive velocity information from the object’s 6-DOF pose at each timestep. This 6-DOF pose information is crucial for the network to learn the rotational dynamics of the objects, in addition to their translational motion. Additionally, we also add features to distinguish between different types of object nodes $\mathcal{V}^{O}$. Our simulation includes kinematic objects (which are controlled) and stationary objects (which remain fixed). To represent this, we use a one-hot encoding to indicate its type. This one-hot encoding is also extended to the mesh nodes $\mathcal{V}^{M}$ since each mesh node belongs to a specific object.
@@
Edge features For the edge sets $\mathcal{E}^{L}$ (liquid-liquid), $\mathcal{E}^{MO}$ (mesh-to-object), and $\mathcal{E}^{OM}$ (object-to-mesh), we adopt the relative encoding technique [sanchez2020learning]. The edge features consist of the relative displacement vector between two connected nodes and the scalar distance between them. The calculation of these relative features differs slightly depending on the edge type. For $\mathcal{E}^{L}$, the relative position and distance are computed between the positions of the two connected liquid particles. For $\mathcal{E}^{MO}$ and $\mathcal{E}^{OM}$ edges, we calculate them based on the mesh node position and the center of mass position of the corresponding object, derived from the object node’s 6-DOF pose. For the liquid-object edge set $\mathcal{E}^{OL}$, which is critical for modeling particle collisions with surfaces, we use a specialized feature set. For an edge connected between particle $p_{j}$ and object $o_{i}$, the feature vector is defined as $e_{ij}^{ol}=[c_{ij}-p_{j},c_{ij}-p_{o_{i}},\lVert c_{ij}-p_{j}\rVert,\lVert c_{ij}-p_{o_{i}}\rVert]$. Here, $c_{ij}$ represents the closest point on the surface of object $o_{i}$ to the particle $p_{j}$, and $p_{o_{i}}$ represents the center of mass of object $o_{i}$. These features together give information about the spatial relationship between the connected particle and object, using the closest point on the object’s surface $c_{ij}$, as a reference. The resulting input graph is encoded into separate MLPs for each node type and edge type before passing to the processor block.